OpenStudy (luigi0210):

Differentiate: \[\LARGE y=sin(5x)\]

OpenStudy (anonymous):

use chain rule or just 5.cos 5x

OpenStudy (luigi0210):

I would use chain rule right? \[\LARGE y'=(sin(5x)'(5x)'\] Which gives me: \[\LARGE y'=5cos5x\]

OpenStudy (anonymous):

y = sin(5x) Let u=5x du/dx = 5 y = sin(u) dy/du = cos(u) = cos(5x) dy/dx = (dy/du)(du/dx) = cos(5x) * 5 = 5cos(5x)